OpenStudy (anonymous):

A curve has a gradient function of ax+b where a and b are constants. Given that the curve has a minimum value of y of 1 when x=1/3 and a gradient of 4 when x=1, find the equation of the curve.

OpenStudy (anonymous):

@myininaya can you help me with this ?

myininaya (myininaya):

Do you know how to integrate?

OpenStudy (anonymous):

a bit.

myininaya (myininaya):

if you know the the curve has the slope formula ax+b then you can find the general form for a curve with that slope by integrating the slope.

OpenStudy (anonymous):

if you dont mind, can u show me the step by drawing? it make me understand more easily

myininaya (myininaya):

integrate (ax+b)?

OpenStudy (anonymous):

ax^2/2a + bx/b

myininaya (myininaya):

that's now how you integrate ax+b

myininaya (myininaya):

a and b are constants how do you integrate x? x^2/2 to integrate 1? you say what?

myininaya (myininaya):

\[\int\limits_{}^{}(ax+b) dx=a \int\limits_{}^{}x dx+b \int\limits_{}^{}1 dx\]

OpenStudy (anonymous):

oh ok. i understand that. and then?

myininaya (myininaya):

what did you get after integrating (ax+b)?

OpenStudy (anonymous):

hmm.. idk

myininaya (myininaya):

have you ever seen this before: \[\int\limits_{}^{}x^n dx =\frac{x^{n+1}}{n+1}+C \text{ where } n \neq -1 \] \[\int\limits_{}^{}k dx=kx+C \text{ where } k \text{ is a constant }\]

OpenStudy (anonymous):

only the upper one. the second one , nope

myininaya (myininaya):

so you don't know how to integrate 5 but x^1?

myininaya (myininaya):

\[\int\limits_{}^{}f dx=F+C \text{ where F'=f}\] Like do you know how to integrate 5?

myininaya (myininaya):

Like think of a function where it's derivative is 5

OpenStudy (anonymous):

aahh, ok...

myininaya (myininaya):

ok what is it?

myininaya (myininaya):

if you can't integrate 5 then I don't know if I can help you with your problem because I have to go to bed soon. I'm getting you example of how to to integrate a constant so you can integrate the constant in your problem. Another word for integrate is find the antiderivative Can you find the antiderivative of 5?

myininaya (myininaya):

Do you know what function gives you the derivative 5?

OpenStudy (anonymous):

no.

myininaya (myininaya):

Do you know anything about derivatives?

OpenStudy (anonymous):

dy/dx?

myininaya (myininaya):

What can you put in this blank so that the below statement is true. \[\frac{d}{dx} ( \text{ ________________) } =5\]

OpenStudy (anonymous):

5x

myininaya (myininaya):

or even 5x+c where c is a constant \[\frac{d}{dx}(5x+c)=5 =>\int\limits_{}^{} 5 dx=5x+c\] So the derivative of 5x+c is 5 so the antiderivative of 5 is 5x+c

myininaya (myininaya):

do you see how to integrate a constant now?

OpenStudy (anonymous):

yes,i can see it now

myininaya (myininaya):

\[\int\limits_{}^{}(ax+b) dx=?\]

myininaya (myininaya):

remember a and b are just constants

OpenStudy (anonymous):

a∫x^2 +c? ax^2 +c? both wrong?

OpenStudy (anonymous):

a∫x^2/2 + b∫x +c

myininaya (myininaya):

here is another example: \[\int\limits_{}^{}(5x+1)dx=5\frac{x^2}{2}+x+c\] here is another another example: \[\int\limits_{}^{} (\pi x-7)dx=\pi \frac{x^2}{2}-7x+c\] now you try again to integrate (ax+b)

myininaya (myininaya):

or did you mean to say earlier it was ax^2/2+bx+c?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok that is what our curve looks like y=ax^2+bx+c has the slope ax+b great now we just need to find a and b and c

myininaya (myininaya):

it tells you what point is on the curve which is what?

OpenStudy (anonymous):

minimum point?

myininaya (myininaya):

yep and what is that point

OpenStudy (anonymous):

idk, maybe (1/3, 1)

myininaya (myininaya):

why is it maybe that is what it says so you can find a relationship between a,b, and c using y=ax^2/2+bx+c using the fact you know (1/3,1) is on the curve.

myininaya (myininaya):

you can replace x with 1/3 and y=1 we will come back in use whatever equation you get from that.

myininaya (myininaya):

you are also given the gradient is ax+b and they tell you some things about the gradient they tell you the gradient is 4 when x=1 so what does that tell you about a relationship just between a and b?

OpenStudy (anonymous):

a is gradient. b is y intercept. ?

myininaya (myininaya):

no it says the gradient is ax+b

myininaya (myininaya):

that is how we were about to find the general form for the curve which is y=ax^2/2+bx+c

myininaya (myininaya):

I want you to tell me a relationship between a and b given the gradient is 4 when x=1

myininaya (myininaya):

what equation does that give you in terms of a and b?

myininaya (myininaya):

gradient(x)=ax+b you are given gradient(1)=4 use that to give me a relation between a and b.

myininaya (myininaya):

you can do this replace gradient(x) with 4 and replace x with 1

OpenStudy (anonymous):

m=ax+b m=4 4=a(1)+b 4=a+b 4-b=a

myininaya (myininaya):

great so we have 4=a+b or you can express it as 4-b=a now you are also given something else about the gradient it says you have a min at x=1/3 what does that mean the gradient is at x=1/3?

myininaya (myininaya):